OpenStudy (anonymous):

Can someone please help - For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 −x − y = −7 I think It's y + 7 is that right?

OpenStudy (anonymous):

it should be x = something. So solve for x in your second equation, -x - y = -7 -x = y - 7 x = -y +7 You had it, but missed the -1 going to the y as well.

OpenStudy (anonymous):

Thanks!